OpenStudy (anonymous):

solve a^2=25 the solution is a=??

OpenStudy (anonymous):

+-5

OpenStudy (anonymous):

how can u have a positive negative 5

OpenStudy (anonymous):

Sheg meant +5 and a -5 would be the answers to this problem

OpenStudy (anonymous):

\[\pm5\]

OpenStudy (anonymous):

oooh lol

OpenStudy (anonymous):

He means positive OR negative 5. Positive 5 squared is positive 25, negative 5 squared is also positive 25.

OpenStudy (anonymous):

because a^2 = 25 a = 25^(1/2) a = + 5 or -5

OpenStudy (anonymous):

(+5)*(+5) = 25 AND (-5)*(-5) = 25